OpenStudy (anonymous):

√9x+55=x+5 Please give me exact answer.

OpenStudy (anonymous):

The answer I keep getting is wrong.

OpenStudy (anonymous):

is it sq root of 9x or just the 9

OpenStudy (karatechopper):

-25 is x is the check meaining radical if so then tht is the answer i just gave u

OpenStudy (anonymous):

sq root

OpenStudy (anonymous):

if its sq root of just the 9.. x= -50 / 2

OpenStudy (karatechopper):

i have the answer already posted for u take my answer

OpenStudy (anonymous):

\[\sqrt{9x+55}=x+5\]

OpenStudy (karatechopper):

u want the steps?

OpenStudy (anonymous):

is it -50/2 is that a sqrt

OpenStudy (karatechopper):

no just do tht equation and then tht is your answer

OpenStudy (anonymous):

-25

OpenStudy (karatechopper):

CORRECT!

OpenStudy (anonymous):

it says that it is wrong

OpenStudy (anonymous):

karatechopper correct same thing -25 is -50/2

OpenStudy (anonymous):

right the answer you gave me was wrong the right answer was 5

OpenStudy (anonymous):

look at this one for me please. \[\sqrt{5x+29}=x+3\]

OpenStudy (anonymous):

5x+29=x ^2 + 6x + 9

OpenStudy (anonymous):

is that the answer.

OpenStudy (anonymous):

half way

OpenStudy (anonymous):

really need the full final answer please

OpenStudy (anonymous):

\[x^{2} + x = 20\]

OpenStudy (anonymous):

x=20

OpenStudy (anonymous):

i make x = 4

OpenStudy (anonymous):

it says that is wrong

OpenStudy (anonymous):

ok. sorry we fail.

OpenStudy (anonymous):

it is cool. just really need the answer.

OpenStudy (anonymous):

try again.

OpenStudy (anonymous):

√ (5x+29) = x+3 i still make it 4. Buts just with pen and paper x is 20 you say. now checking if u know thats the answer why ask us

OpenStudy (anonymous):

ahh \[x = x ^{2}+x\]

OpenStudy (anonymous):

makes x=0. cant be right. lol. i give up. g luck